# Early release of low-frequency atlas of continuous gravitational waves

### v1

@@ Abstract
We present the public release of the low-frequency atlas of continuous gravitational waves, covering signals with frequencies from 20 Hz to 200 Hz and frequency derivatives from $-5\times 10^{-11}$ to $5\times 10^{-11}$ Hz/s. Compared to the previous atlas releases, this version demonstrates significant improvements in sensitivity and sky resolution. In the most sensitive region even the worst-case upper limits on gravitational wave strain are below $10^{-25}$. The atlas data is being released ahead of the completion of the full follow-up analysis.
@@ I Introduction
The first continuous gravitational wave atlas was released several years ago [2] and it covered signals with frequencies between 500-1000 Hz and frequency derivatives from $-5\times 10^{-11}$ to $5\times 10^{-11}$ Hz/s. Since then, we produced new atlases that progressively expanded parameter space coverage [2, 3, 4].
@@ II Signal model
where $I_{zz}=10^{38}$ kg m2 is the moment of inertia of the star with respect to the principal axis aligned with the rotation axis. Gravitational waves carry away energy, gradually slowing down the pulsar’s rotation. Most known radio pulsars have frequency derivatives below with ±$\pm 5\times 10^{-11}$ Hz/s.
To compute the atlas data we assume a source which emits gravitational waves with the following frequency evolution in the Solar System baricenter
@@
Higher order frequency derivatives are not explicitly searched over, instead relying on intrinsic robustness of a search with coherence lengths of 2-days and using one year of data.
@@ III Atlas data
3 $108.85716$1.46e-17 34.1 126.7 189.4 0.031 Yes
5 $52.80832$4.03e-18 259.9 135.2 364.6 -0.005 Yes
6 $145.33976$6.73e-09 8.2 19.8 638.0 -30.505 No
8 $189.96498$8.65e-09 8.3 58.8 215.0 0.832 No
10 $26.33144$8.5e-11 51.2 86.4 120.5 -23.734 No
11 $31.42470$5.07e-13 41.7 245.4 87.4 -10.257 Yes
12 $37.70750$6.25e-09 9.7 106.4 133.3 47.890 No
Table 1: This table shows the frequency parameters of the hardware-injected continuous wave signals and how they appear in the atlas data. We show all the hardware injections within 20-200 Hz range, including those outside of our search space, as indicated by the “In” column. We show our upper limit UL at the injection sky position and polarization parameter values, and how it compares with the injection amplitude $h_{0}$: if the upper limits are correct, UL/$h_{0}$ has to be $>100$%. $h_{0}$/LL shows the ratio of signal strain to the lock level, and we expect to detect signals when this ratio is $>100\%$. This is the case for the first two hardware-injected signals in range, but not for the third, as it can be seen from the large mismatch difference $\Delta f$ between the recovered frequency and the signal frequency. For detectable signals in our parameter space we expect the $\Delta f$ to be within $0.5$ mHz. We use the reference time (GPS epoch) $t_{0}=1253800000$ s (2019 Sep 29 13:46:22 UTC). The script spatial_index_example2.R provided with the atlas shows how to obtain this data from the atlas.
@@ IV Population average proxies
The location-specific polarization-average lock-level proxy is defined as the signal strain level at which the frequency and polarization parameters of the SNR peak in the correspoding Voronoi domain are correctly reported in the atlas for 95% of signals. The primary intended use for the proxy is to assess when a limited parameter-space follow-up of SNR-peak parameters has merit, or when the entire Voronoi domain needs to be reanalyzed with a more sensitive search in the given frequency band.
Our location-specific polarization-average lock-level proxies are computed for each atlas entry from an estimate of the background noise and are insensitive to the presence of plausible strength signals. In our injection study 99% of injections changed the lock level by less than 1%. We validate these proxies using test signals distributed uniformly on the sky and with frequencies in the 100-200 Hz band. For each test signal, we identify the nearest point in the atlas and compare the SNR peak parameters for that point to those of the test signal. Figure 2 shows the percentage of signals with SNR peak frequencies within $0.5$ mHz of the actual value, as well as the percentage of signals with both frequency and polarization correctly recovered, as a function of signal strength. We verify that when the signal strength is at the lock-level, the recovered fraction of signals across a population of signals with arbitrary polarization is 95%. Figure 2 also shows that frequencies are recovered most accurately, while polarization recovery is less precise. This is because polarization reconstruction relies heavily on amplitude modulation, introducing significant degeneracy, especially for nearly circularly polarized signals (Figures 3 and 4).
Location-specific lock-level proxies can be used to estimate the sensitivity of our search for a population of signals coming from any direction in the sky, and this can in turn be used to generate an estimate of the so-called population-average upper limits over the whole sky, in the absence of signals. This is useful because most all-sky searches [17, 16] report as population-average upper limits. These are conceptually simple to implement by simulating search results from populations of signals and comparing with the observation results, until the signal amplitude that yields the 95% quantile is identified. However, practical challenges may arise, and the computational cost of these estimates is generally not trivial. The Falcon searches, like the one presented here, and its predecessor the Powerflux search, have used a different approach, establishing upper limits for fixed polarizations using the universtal statistics [15], which do not suffer from the same limitations. However, this can make comparison with the results of other pipelines more difficult.
@@
Figure 2: Signal recovery versus the ratio of signal strain to lock-level estimated on data with the test signals. The red line marks the $95$% threshold. The blue crosses show the percentage of points where the SNR peak frequency was withing $0.5$ mHz of the injected value.Given the $50$ mHz frequency band there is a 1% chance of false coincidence. The magenta circles show the percentage of points where $\iota$ was within $0.7/\sin(\iota)$ of injected value and $\psi$ was also within $0.7$ of injected value modulo $\pi/2$. Such wide tolerances are due to shoulders in $\iota$ and $\psi$ recovery (see Figures 3 and 4).
The all-sky population average upper limit proxy shown in Figure 1 are the 95% quantile of the location-specific population-average upper limit proxies, resulting in a 90% confidence level for the curve shown in Figure 1 which is meant to be used for pipeline comparison purposes only.
@@
Figure 4: Recovery of the polarization angle $\psi$ from the SNR peaks. We plot only points corresponding to injected strain at or above the polarization-average lock-level. The left plot restricts data to nearly linearly polarized signals. The right plot includes all points, but multiplies the difference between injected and recovered $\psi$ values (modulo $\pi/2$) by $\sin(\iota)$. This way the mismatch in $\psi$ is deweighted as $\iota$ approaches circular polarizations 00 or $\pi$ (the “poles” of the polarization sphere).
@@ V Conclusions
We have conducted a wide-band search of the full O3 data covering frequencies from 20 to 200 Hz. The results have been compiled into an atlas, which we release while the final analysis is ongoing. We introduce population average lock-level proxies for every sky location in the atlas data. The atlas includes several examples, written in R, that can be a starting point for future searches.

### v2

@@ Abstract
We present the public release of the low-frequency atlas of continuous gravitational waves, covering signals with frequencies from 20 Hz to 200 Hz and frequency derivatives from $5\times 10^{-11}$ to $5\times 10^{-11}$ Hz/s. Compared to the previous atlas releases, this version demonstrates significant improvements in sensitivity and sky resolution. In the most sensitive region even the worst-case upper limits on gravitational wave strain are below $10^{-25}$. The atlas data is being released ahead of the completion of the full follow-up analysis.
@@ I Introduction
The first continuous gravitational wave atlas was released several years ago [2] and it covered signals with frequencies between 500-1000 Hz and frequency derivatives from $5\times 10^{-11}$ to $5\times 10^{-11}$ Hz/s. Since then, we produced new atlases that progressively expanded parameter space coverage [2, 3, 4].
@@ II Signal model
where $I_{zz}=10^{38}$ kg m2 is the moment of inertia of the star with respect to the principal axis aligned with the rotation axis. Gravitational waves carry away energy, gradually slowing down the pulsar’s rotation. Most known radio pulsars have frequency derivatives below $\pm 5\times 10^{-11}$ Hz/s.
To compute the atlas data we assume a source which emits gravitational waves with the following frequency evolution in the Solar System barycenter
@@
Higher order frequency derivatives are not explicitly searched over, instead relying on intrinsic robustness of a search with coherence lengths of 2 days and using one year of data.
@@ III Atlas data
3 $108.85716$ -1.46e-17 34.1 126.7 189.4 0.031 Yes
5 $52.80832$ -4.03e-18 259.9 135.2 364.6 -0.005 Yes
6 $145.33976$ -6.73e-09 8.2 19.8 638.0 -30.505 No
8 $189.96498$ -8.65e-09 8.3 58.8 215.0 0.832 No
10 $26.33144$ -8.5e-11 51.2 86.4 120.5 -23.734 No
11 $31.42470$ -5.07e-13 41.7 245.4 87.4 -10.257 Yes
12 $37.70750$ -6.25e-09 9.7 106.4 133.3 47.890 No
[…]
@@ IV Population average proxies
The location-specific polarization-average lock-level proxy is defined as the signal strain level at which the frequency and polarization parameters of the SNR peak in the corresponding Voronoi domain (sky region around a sky-grid point and 50 mHz frequency band) are correctly reported in the atlas for 95% of signals. The primary intended use for the proxy is to assess when a limited parameter-space follow-up of SNR-peak parameters has merit, or when the entire Voronoi domain needs to be reanalyzed with a more sensitive search in the given frequency band.
Our location-specific polarization-average lock-level proxies are computed for each atlas entry from an estimate of the background noise and are insensitive to the presence of plausible strength signals. In our injection study 99% of the test signals (of Figure 2) change the lock level by less than 1%.
We validate this proxy using test signals distributed uniformly on the sky and with frequencies in the 100-200 Hz band. For each test signal, we consider the parameters of the SNR peak at the sky-point closest to the position of the test signal and in the relevant 50 mHz frequency band. Figure 2 shows the percentage of signals with SNR peak frequency within $0.5$ mHz of the actual value, as well as the percentage of signals with both frequency and polarization correctly recovered, as a function of signal strength. We verify that when the signal strength is at the lock-level, the recovered fraction of signals across a population of signals with arbitrary polarization is at the 95% level. Figure 2 also shows that frequencies are recovered most accurately, while polarization recovery is less precise. This is because polarization reconstruction relies heavily on amplitude modulation, introducing significant degeneracy, especially for nearly circularly polarized signals (Figures 3 and 4).
Location-specific lock-level proxies can be used to estimate the sensitivity of our search for a population of signals coming from any direction in the sky, and this can in turn be used to generate an estimate of the so-called population-average upper limits over the whole sky, in the absence of signals. This is useful because most all-sky searches [19, 16, 17, 18] report results population-average upper limits. These are conceptually simple to implement by simulating search results from populations of signals and comparing with the observation results, until the signal amplitude that yields the 95% quantile is identified. However, practical challenges may arise, and the computational cost of these estimates is generally not trivial. The Falcon searches, like the one presented here, and its predecessor the Powerflux search, have used a different approach, establishing upper limits for fixed polarizations using the universal statistics [15], which do not suffer from the same limitations. However, this can make comparison with the results of other pipelines more difficult.
@@
Figure 2: Signal recovery versus the ratio of signal strain to lock-level estimated on data with the test signals. The red line marks the 9595% threshold. The blue crosses show the percentage of points where the SNR peak frequency was withing $0.2$ mHz of the injected value. Given the $50$ mHz frequency band there is a 0.4% chance of false coincidence. The magenta circles show the percentage of points where $\iota$ was within $0.7/\sin(\iota)$ of injected value and $\psi$ was also within $0.7$ of injected value modulo $\pi/2$. Such wide tolerances are due to shoulders in $\iota$ and $\psi$ recovery (see Figures 3 and 4).
@@
Figure 4: Recovery of the polarization angle $\psi$ from the SNR peaks. We plot only points corresponding to injected strain at or above the polarization-average lock-level. The left plot restricts data to nearly linearly polarized signals. The right plot includes all points, but multiplies the difference between injected and recovered $\psi$ values (modulo $\pi/2$) by $\sin(\iota)$. This way the mismatch in $\psi$ is deweighted as $\iota$ approaches circular polarizations 0 or $\pi$ (the “poles” of the polarization sphere).
Band start Band stop SNR Frequency Frequency
Hz Hz all-sky lock, % & pol lock, %
100.003 100.054 98.2 64.0 61.0
100.053 100.104 17.8 99.0 96.5
100.103 100.154 10.6 99.5 98.0
100.153 100.204 10.3 99.5 94.0
100.203 100.254 10.2 99.0 98.0
100.253 100.304 12.3 99.5 97.5
100.303 100.354 11.3 99.5 98.0
100.353 100.404 10.9 99.5 97.0
100.403 100.454 10.3 100.0 97.5
100.453 100.504 10.5 97.0 92.5
100.503 100.554 10.4 98.0 91.5
100.553 100.604 10.3 99.0 97.0
100.603 100.654 10.4 99.5 95.5
100.653 100.704 10.6 99.0 97.0
100.703 100.754 10.5 100.0 95.0
100.753 100.804 10.6 99.5 95.5
100.803 100.854 13.5 97.5 93.0
100.853 100.904 10.5 99.5 95.5
100.903 100.954 10.4 98.0 96.0
100.953 101.004 10.7 100.0 96.0
Table 2: This table shows how signals are recovered in the 100-101 Hz frequency range and it is an illustrative spot-check of the all-sky population average upper limit estimate. In each band we consider 200 test-signals with strain equal to the all-sky population-average upper limit estimate and the other parameters distributed randomly in the search parameter space. We add each signal to the data and carry out the search. Columns 4 and 5 show the percentage of test-signals for which the frequency (column 4) and frequency+polarization (column 5) are correctly recovered in the test signal parameters Voronoi domain. For signals with arbitrary sky position we expect both these numbers, and in particular that of column 5, to be at least 90% , if the all-sky population-average upper limit estimate is correct. Indeed we see that in most frequency bands this is the case. In the band containing a strong disturbance, such as the first band of the table, the SNR peak is due to the disturbance and not representative of a signal, so “signal-lock” at the proxy level is not achieved as effectively. Column 3 shows the highest SNR value from each band in the atlas results. Large value indicate a loud signal, astrophysical or not.
Figure 5: This figure shows how SNR varies over sky position in the band 100.003-100.054 Hz. We observe strong contamination from instrumental artifacts near 100 Hz. The yellow plus signs denote locations of injections with $|\cos(\iota)|<0.5$ with frequency lock. The injections group in areas with smaller SNR away from contaminated areas.
We have performed spot-checks of the all-sky population average upper limit estimates, the results of which are shown in Table 2. Each frequency band was tested with 200 simulated signals having strain equal to all-sky population average proxy and the rest of parameters distributed randomly. We observe that lock rates are above the design value of 90% for all bands except 100.003-100.054 Hz, which is highly contaminated by detector artifacts (Figure 5).
Figure 5 shows that nearly linearly-polarized injections fail to achieve frequency lock in contaminated regions with high-SNR from detector artifacts. This is expected behaviour because the lock-level is a measure of background noise and is constructed to be insensitive to any signals in the band - whether astrophysical or technical.
More generally, there is a natural variation in detector sensitivity with sky position. Thus the rate of test-signal recovery based on any population average proxy, regardless of how it is constructed, will vary with the position of the source in the sky. This again highlights the dependence of population average proxy on the population it was designed for, in contrast to frequentist upper limits designed with universal statistics [15], which apply to any subset of parameter space.
@@ V Conclusions
We have conducted a wide-band search of the full O3 data covering frequencies from 20 to 200 Hz, which is the most sensitive all-sky search to date. The results have been compiled into an atlas, which we release while the final analysis is ongoing. We introduce population average lock-level proxies for every sky location in the atlas data. The atlas includes several examples, written in R, that can be a starting point for future searches.
Acknowledgements.
The authors thank the scientists, engineers and technicians of LIGO, whose hard work and dedication produced the data that made this search possible. The search was performed on the ATLAS cluster at AEI Hannover. We thank Bruce Allen, Carsten Aulbert and Henning Fehrmann for their support. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gw-openscience.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA.